# A localized consensus-based sampling algorithm

### v2

@@ 7 Numerical results
In this section, we apply the localized CBS algorithm to a variety of problems, illustrating its properties. For both localized CBS and alternative methods, we will use Euler–Maruyama time discretizations. Unless mentioned otherwise, localized CBS uses timestep ${\Delta\kern-0.8ptt}=0.01$ and preconditioner $\mathcal{C}(\rho_{t})$, an unweighted covariance. The default random-batch parameter is $\nu=1$, such that a particle interacts with all other particles. In addition, except in subsection 7.3, we obtain samples of all considered methods by running the algorithm with the specified parameters $16$ times, and combining the particle positions at all time steps within the final quarter of the time interval in each run. Samples are plotted with the density function from the Julia package StatsPlots, which performs kernel density estimation (KDE) to smooth the samples into an approximate distribution. When the dimension $d$ is larger than one, the plots show the marginal distribution in the first dimension unless specified otherwise. The initial distribution for all sampling methods will be a zero-centered Gaussian $\mathcal{N}(0,\Sigma_{0})$; by default, ${\Sigma_{0}=I/2}$. Our code is located at
@@ 7.1 Gaussian distributions
and use $(\beta,\kappa)=(5.0,0.01)$, $J=500$ particles, and $N=200$ timesteps. Figure 1 shows the resulting samples with $\gamma\in\{0.5,\bar{\gamma},1.0,1.5\}$, where ${\bar{\gamma}\approx 0.843}$ is the value given by Eq. 38—since we use our default preconditioner $\mathcal{C}(\rho_{t})$, this should ensure correct sampling for Gaussians. When $\gamma<\bar{\gamma}$, there is insufficient drift towards the weighted mean and the samples are too spread out. When $\gamma>\bar{\gamma}$, the opposite effect occurs. As predicted by the theory, the distribution is sampled correctly with $\gamma=\bar{\gamma}$.
@@ 7.2 Multimodal distributions
For this potential and with ${d\in\{1,10\}}$, we compare localized CBS (with $\beta=10$ and $\kappa=0.03$) to CBS (with $\alpha=10$ and ${\Delta\kern-0.8ptt}=0.01$), polarized CBS (with $\alpha=10$, $\lambda=0.005$ when $d=1$ or $\lambda=0.1$ when $d=10$, $D=I$, and ${{\Delta\kern-0.8ptt}=0.01}$), and localized ALDI (with $\lambda=0.02$ when $d=1$ or $\lambda=0.4$ when $d=10$, $D=I$, and ${\Delta\kern-0.8ptt}=0.02$). All tests use $J=200$ particles and $N=1000$ timesteps. For localized CBS, we additionally compare $\nu=1$ to $\nu=0.5$. Figure 2 shows the resulting samples. Localized CBS accurately finds the distribution when $d=1$; when $d=10$, we get a good approximation only with random-batch interaction turned on ($\nu=0.5$). Localized ALDI and polarized CBS each give relatively close approximations, but are unable to exactly fit $\pi$. Note also that for these two methods, unlike for localized CBS, we needed to hand-tune the parameters for each $d$. Classical CBS finds the multiple modes only when $d=10$, and fails to approximate the density in between. [^7] Note that the particles in CBS always follow a Gaussian distribution; in our experiment, CBS captures multiple modes as the results from $16$ runs (which may find different modes) are aggregated.
@@
where $u$ is $(d=2)$-dimensional and the square again applies elementwise. The matrix $\Lambda=\mathrm{diag}(1,10^{4})$ changes the scale of the second parameter. Suppose that we may not know this scaling before solving the problem. Without knowing the form of the potential Eq. 73, we assume that the relative scaling of the two parameters is given by a scaling matrix $\tilde{\Lambda}$ that can either be the correct $\tilde{\Lambda}=\Lambda$ or the incorrect $\tilde{\Lambda}=I$. We use this assumption to choose an initial covariance $\Sigma_{0}=\tilde{\Lambda}^{-1}/2$. Furthermore,
@@
We use $(\beta,\kappa)=(10,0.03)$ for localized CBS, $\lambda=0.02$ and ${\Delta\kern-0.8ptt}=0.05$ for localized ALDI, and $\lambda=0.005$ for polarized CBS. Figure 3 shows the values sampled by these algorithms for both $\tilde{\Lambda}$ cases. Polarized CBS and localized ALDI perform as expected when the scaling is guessed correctly, but generate clearly incorrect samples when it is not. This is because they depend on $\tilde{\Lambda}$ not only for their initial condition, but also for weighting the distance between particles throughout the algorithm. Localized CBS, on the other hand, proves itself able to recover from this poor initial guess and produces good—and almost identical—samples for both $\tilde{\Lambda}$s.
Figure 3: Marginalized samples for a 2-dimensional poorly scaled multimodal target distribution. In the left figure, the two red lines are visually indistinguishable; in the middle figure, the line for $\tilde{\Lambda}=I$ coincides with the horizontal axis as the samples diverge.
@@
Next, we illustrate the importance of the correction term $\nabla_{u}\cdot\mathbf{C}_{U,t}$ in the localized CBS dynamics Eq. 29 on the sampling problem Eq. 74. We consider the preconditioner $\mathcal{C}_{0,\lambda\mathcal{C}(\rho_{t})}(\rho_{t};U_{t})$ with $\lambda=0.5$ and compare the normal localized CBS algorithm to a modified version where the correction term is omitted. We also apply polarized CBS, which does not have a similar correction term, with $\alpha=10$ and $\lambda\in\{0.001,0.002,0.005\}$. We use $N=1000$ timesteps, $J=200$ particles, and an initial covariance $\Sigma_{0}=2I$ for all samplers. Figure 5 illustrates that only localized CBS with the correction term manages to accurately retrieve $\pi$. As before, polarized CBS only samples from a distribution approximating $\pi$.
@@ 7.3 A tent-shaped distribution
a tent-shaped distribution with finite support. We will answer the question why, even though $\beta\rightarrow\infty$ and $\kappa\rightarrow 0$ seem desirable, it may still make sense to choose moderate values for both parameters. In Figure 6, we show the Wasserstein-2 distance between the exact distribution $\pi$ and the empirical distribution of localized CBS ensembles with a given number $J$ of particles. This is compared for several $(\beta,\kappa)$ combinations. In this subsection, we find the empirical distribution for any $(\beta,\kappa,J)$ by simulating localized CBS $480$ times and, for each simulation, using the positions of $50$ randomly chosen particles at the final time step as samples.
Figure 6: Wasserstein-$2$ error between the localized CBS samples and the target distribution Eq. 75
@@
When $J$ keeps increasing, the errors in Figure 6 become slightly larger again. This could be caused by a combination of two factors: (i) the Moreau envelope step in section 3 approximates $\pi$ by a distribution with longer tails and (ii) when $J$ is small, localized CBS tends to undersample the tails. While the first factor is a source of error that is present for all $J$, the second factor may partially offset it for small $J$, resulting in an error smaller than that for $J\rightarrow\infty$.
@@ 7.4 A Darcy flow inverse problem
Consider a two-dimensional spatial domain $\Omega\coloneqq[0,1]^{2}$. The Darcy flow inverse problem is to recover parameters $u$ of the permeability field $a(x,u)$ from noisy measurements of the pressure field $p(x)$ satisfying
@@
The permeability $a(x,u)$ is modeled as a log-normal random field with mean zero and covariance operator $(-\Delta+\tau^{2})^{-s}$ of $\log a$, where $-\Delta$ represents the Laplacian with homogeneous Neumann boundary conditions. This corresponds to a standard normal prior on the parameters $u_{k}$ in the Karhunen–Loève (KL) expansion
$\log a(x,u)=\sum_{k\in\mathbb{N}^{2}\setminus\{(0,0)\}}u_{k}\sqrt{\lambda_{k}}\,\phi_{k}(x),$ (77)
with eigenpairs given by $\lambda_{k}=(\pi^{2}\lVert k\rVert^{2}+\tau^{2})^{-s}$ and $\phi_{k}=c_{k}\cos(\pi k_{1}x_{1})\cos(\pi k_{2}x_{2})$, where ${c_{k}=\sqrt{2}}$ if $k_{1}k_{2}=0$ and $c_{k}=2$ otherwise (see, e.g., [27]). We choose parameters $(\tau,s)=(3,2)$ and truncate the expansion Eq. 77 after the $d$ terms with largest eigenvalues $\lambda_{k}$. The coefficients $u_{k}$ of these terms are the unknown parameters $u$ with prior $\pi_{\mathrm{prior}}=\mathcal{N}(0,\Gamma_{0})$, with $\Gamma_{0}=I$, in the Bayesian inverse problem Eq. 2. The data $y$ consists of measurements of $p$, taken at $49$ equispaced points and subject to noise known to be distributed according to $\pi_{\mathrm{noise}}=\mathcal{N}(0,\Gamma)$ with $\Gamma=10^{-4}I$. The model $G$ solves Eq. 76 with central finite differences on a grid with stepsize $h=2^{-5}$. An observation is generated by sampling $\{u_{k}\}_{k}$ from the prior, solving Eq. 76 with stepsize $h=2^{-9}$, and perturbing that solution with a random sample from $\pi_{\mathrm{noise}}$. As described in subsection 1.1, this Bayesian inverse problem corresponds to Eq. 1 with
@@
Figure 7: Random-walk Markov chain Monte Carlo and localized CBS approximations to the distributions of the first $8$ KL terms in the $d$-dimensional Darcy flow problem, with $d\in\{8,16,32\}$. The $d=8$ plot labels the modes in order of decreasing eigenvalue $\lambda_{k}$.
We solve this inverse problem with random-walk MCMC (as a reference solution) and localized CBS. We use $N=400$ timesteps of length ${\Delta\kern-0.8ptt}=0.03$${J=1000}$ particles, and a large initial covariance $\Sigma_{0}=9I$ (similarly to [6]). Since $\Gamma$ is small, we expect the posterior to be close to Gaussian, which is also confirmed by the MCMC reference solution. Hence, we choose moderate parameters $(\beta,\kappa)=(3,0.2)$. Figure 7 compares the MCMC samples to the localized CBS samples. It shows the marginalized distributions for the first 8 modes, for the number of dimensions $d\in\{8,16,32\}$. The modes are successfully located for all $d$ values. For each $d$, the MCMC modes are recovered accurately by localized CBS. When $d=8$ or $d=16$, the uncertainty around the mode is captured well; when $d=32$, it is slightly overestimated.
@@ 8 Conclusions
We proposed a novel interacting-particle method for sampling from non-Gaussian distributions, based on an alternative interpretation of consensus-based sampling (CBS) as an approximation to interacting Langevin diffusions. In the mean-field limit, our localized CBS method was shown to have desirable properties such as affine-invariance and unbiased sampling from Gaussian distributions.
Numerical experiments showed that localized CBS deals well with poorly scaled sampling problems, even when this scaling is unknown in advance. It is also able to accurately identify multiple modes in problems with low or moderate dimension.
Localized CBS has several parameters, including a preconditioner that can be chosen very broadly. In our experiments, we have focused on the simple preconditioner $\mathcal{C}(\rho_{t})$—an unweighted covariance of the SDE’s law at time $t$—and briefly explored a more complex space-dependent preconditioner. Future research could focus on studying the choice of the preconditioner and other parameters in more detail. In addition, it would be very valuable to obtain convergence results for localized CBS with a finite number of particles. These results could also elucidate the precise role played by random-batch interaction, which appears critical to the performance of localized CBS in some multimodal higher-dimensional problems.
More broadly, a detailed study of the influence of timestepping schemes on the efficiency and accuracy of samplers based on interacting particles is likely to make these methods more practical as a whole. In this context, the adaptive timestepping scheme developed in [32] for ensemble Kalman inversion—and used in [19] for ensemble Kalman sampling—is a promising example.
@@ Appendix A Various technical proofs
In this appendix, we give proofs for some results used in the main text.
Proof of Eqs. 26, 34 and 35.
Consider means $(m,m_{t})$ and covariances $(\Sigma,\Sigma_{t})$ such that ${\pi\sim\mathcal{N}(m,\Sigma)}$ and $\rho_{t}\sim\mathcal{N}(m_{t},\Sigma_{t})$. We prove Eq. 26 by computing
@@
which is the mean of the product of two Gaussians, $\mathcal{N}(m,\Sigma/\alpha)$ and $\mathcal{N}(m_{t},\Sigma_{t})$. We use the formula found in, e.g., [45, §8.1.8] to compute $\mu_{\alpha}(\rho_{t})$ and, similarly, $\mathcal{C}_{\alpha}(\rho_{t})$:
$\mu_{\alpha}(\rho_{t})=(\alpha\Sigma^{-1}+\Sigma_{t}^{-1})^{-1}(\alpha\Sigma^{-1}m+\Sigma_{t}^{-1}m_{t})\;\;\text{and}\;\;\mathcal{C}_{\alpha}(\rho_{t})=(\alpha\Sigma^{-1}+\Sigma_{t}^{-1})^{-1}.$ (80)
@@
Equations 34a and 34b follow directly from the definitions. Equations 34c, 34d and 35 are derived similarly to $\mu_{\beta,\kappa\mathcal{C}_{\alpha}(\rho_{t})/\beta}(\rho_{t};U_{t})$ above. ∎
@@
In [47, Lemma 3.11], the divergence of a weighted covariance with general weight function is given:
$\nabla_{u^{i}}\cdot\widehat{\mathbf{C}}=\omega^{ii}\,(d+1)(U^{i}-\bar{U}^{(i)})+\sum\nolimits_{j=1}^{J}\left(U_{c}^{(i),j}(U_{c}^{(i),j})^{T}-\bar{U}^{(i)}(\bar{U}^{(i)})^{T}\right)\nabla_{u^{i}}\omega^{ij}.$ (81)
@@
$\quad=\frac{\omega^{ij}}{\lambda}\Bigl{(}\bigl{[}\widehat{\mathcal{C}}^{-1}(U^{j}-U^{i})+\frac{X^{ij}}{2}\bigr{]}+\bigl{[}\widehat{\mathcal{C}}^{-1}U_{c}^{(i),i}-\sum\nolimits_{k=1}^{J}\omega^{ik}\frac{X^{ik}}{2}\bigr{]}\Bigr{)}$
$\quad\qquad+(\delta^{ij}-\omega^{ij})\alpha\omega^{ii}\frac{\nabla_{u^{i}}\widehat{\pi}(U^{i})}{\widehat{\pi}(U^{i})}$
@@
where $X^{ik}$ is a vector whose $\ell$th element is $(U^{i}-U^{k})^{T}\widehat{\mathcal{C}}^{-1}\,(\partial_{u^{i}_{\ell}}\widehat{\mathcal{C}})\,\widehat{\mathcal{C}}^{-1}(U^{i}-U^{k})$, with $\partial_{u^{i}_{\ell}}\widehat{\mathcal{C}}=J^{-1}\bigl{(}U_{c}^{i}e_{\ell}^{T}+e_{\ell}(U_{c}^{i})^{T}\bigr{)}$. It can be verified that
@@
$B_{1}$ $\coloneqq\sum\nolimits_{j=1}^{J}U_{c}^{(i),j}(U_{c}^{(i),j})^{T}\omega^{ij}\bigl{[}\widehat{\mathcal{C}}^{-1}U_{c}^{(i),j}\bigr{]}$
$=\bm{U}_{c}^{(i)}\bigl{(}\omega^{i}\odot\mathrm{diag}((\bm{U}_{c}^{(i)})^{T}\widehat{\mathcal{C}}^{-1}\bm{U}_{c}^{(i)})\bigr{)},$

### v3

@@ 7 Numerical results
This section applies the localized CBS algorithm to a variety of problems, illustrating its properties. We compare our method to several existing alternatives: CBS (to illustrate that it is not suited for non-Gaussian problems) on the one hand and polarized CBS and localized ALDI on the other (to compare our algorithm to the two most similar options). Other algorithms are often challenging to compare to in a fair way: they either use gradients, are less parallelizable, or produce fewer samples per evaluation of $V$.
For both localized CBS and alternative methods, we will use the Euler–Maruyama time discretizations discussed in section 6. Unless mentioned otherwise, localized CBS uses timestep ${\Delta\kern-0.8ptt}=0.01$ and preconditioner $\mathcal{C}(\rho_{t})$, an unweighted covariance. The default random-batch parameter is $\nu=1$, such that a particle interacts with all other particles. In addition, except in subsection 7.3, we sample by running the specified algorithm $16$ times and combining the particle positions at all time steps within the final quarter of the time interval in each run. Samples are plotted with the density function from the Julia package StatsPlots, which performs kernel density estimation (KDE) to smooth the samples into an approximate distribution. When the dimension $d$ is larger than one, we plot the marginal distribution in the first dimension unless specified otherwise. The initial distribution for all sampling methods is a zero-centered Gaussian $\mathcal{N}(0,\Sigma_{0})$; by default, $\Sigma_{0}=I/2$}. Our code is located at
@@ 7.1 Gaussian distributions
and use $(\beta,\kappa)=(5.0,0.01)$, $J=500$ particles, and $N=200$ timesteps. Figure 1 shows [^7] In the figures of this section, we abbreviate localized CBS, polarized CBS, and localized ALDI as LCBS, PCBS, and LALDI, respectively. the resulting samples with $\gamma\in\{0.5,\bar{\gamma},1.0,1.5\}$, where $\bar{\gamma}\approx 0.843$} is the value given by Eq. 36—since we use our default preconditioner $\mathcal{C}(\rho_{t})$, this should ensure correct sampling for Gaussians. When $\gamma<\bar{\gamma}$, there is insufficient drift towards the weighted mean and the samples are too spread out. When $\gamma>\bar{\gamma}$, the opposite effect occurs. As predicted by the theory, the distribution is sampled correctly with $\gamma=\bar{\gamma}$.
@@ 7.2 Multimodal distributions
For this potential with $d\in\{1,10\}$}, we compare localized CBS (with $\beta=10$} and $\kappa=0.03$}) to CBS (with $\alpha=10$} and ${\Delta\kern-0.8ptt}=0.01$}), polarized CBS (with $\alpha=10$$\lambda=0.005$} when $d=1$ or $\lambda=0.1$ when d=10d=10D=ID=I, and ${\Delta\kern-0.8ptt}=0.01$}), and localized ALDI (with $\lambda=0.02$ when $d=1$ or $\lambda=0.4$ when $d=10$, $D=I$, and ${\Delta\kern-0.8ptt}=0.02$). All tests use $J=200$ particles and N=1000{N=1000} timesteps. For localized CBS, we additionally compare $\nu=1$ to $\nu=0.5$}. Figure 2 shows the resulting samples. Localized CBS accurately finds the distribution when $d=1$; when d=10{d=10}, we get a good approximation only with random-batch interaction turned on ($\nu=0.5$). Localized ALDI and polarized CBS each give relatively close approximations, but are unable to exactly fit $\pi$. Note also that for these two methods, unlike for localized CBS, we needed to hand-tune the parameters for each $d$. Classical CBS finds the multiple modes only when d=10{d=10}, and fails to approximate the density in between. [^8] Note that the particles in CBS always follow a Gaussian distribution; in our experiment, CBS captures multiple modes as the results from $16$ runs (which may find different modes) are aggregated.
@@
where $d=2$ and the square again applies elementwise. The matrix $\Lambda=\mathrm{diag}(1,10^{4})$} changes the scale of the second parameter. Suppose that we may not know this scaling before solving the problem. Without knowing the form of the potential Eq. 69, we assume that the relative scaling of the two parameters is given by a matrix $\tilde{\Lambda}$ that can either be the correct $\tilde{\Lambda}=\Lambda$ or the incorrect $\tilde{\Lambda}=I$. We use this assumption to choose an initial covariance $\Sigma_{0}=\tilde{\Lambda}^{-1}/2$}. In addition,
@@
We use $(\beta,\kappa)=(10,0.03)$ for localized CBS, $\lambda=0.02$ and ${\Delta\kern-0.8ptt}=0.05$ for localized ALDI, and $\lambda=0.005$} for polarized CBS. Figure 3 shows the samples by these algorithms for both $\tilde{\Lambda}$ cases. Polarized CBS and localized ALDI perform as expected when the scaling is guessed correctly, but generate clearly incorrect samples when it is not. This is because they depend on $\tilde{\Lambda}$ not only for their initial condition, but also for weighting the distance between particles throughout the algorithm. Localized CBS, on the other hand, proves itself able to recover from this poor initial guess and produces good—and almost identical—samples for both $\tilde{\Lambda}$s.
Figure 3: Marginalized samples for the poorly scaled distribution with potential Eq. 69. In the left figure, the two red lines are visually indistinguishable; in the middle figure, one line coincides with the horizontal axis.
@@
Next, we illustrate the importance of the correction term $\nabla_{u}\cdot\mathbf{C}_{U,t}$ in the localized CBS dynamics Eq. 28 on the sampling problem Eq. 70. We consider the preconditioner $\mathcal{C}_{0,\lambda\mathcal{C}(\rho_{t})}(\rho_{t};U_{t})$ with $\lambda=0.5$ and compare the normal localized CBS algorithm to a modified version where the correction term is omitted. We also apply polarized CBS, which does not have a similar correction term, with $\alpha=10$ and $\lambda\in\{0.001,0.002,0.005\}$. We use $N=1000$ timesteps, J=200{J=200} particles, and an initial covariance $\Sigma_{0}=2I$ for all samplers. Figure 5 illustrates that only localized CBS with the correction term manages to accurately retrieve $\pi$. As before, polarized CBS only samples from a distribution approximating $\pi$.
@@ 7.3 A tent-shaped distribution
a tent-shaped distribution with finite support. We will answer the question why, even though choosing $\beta\rightarrow\infty$} and $\kappa\rightarrow 0$ seems desirable, it may still make sense to choose moderate values for both parameters. Figure 6 shows the Wasserstein-2 distance between the exact distribution $\pi$ and the empirical distribution of localized CBS ensembles with a given number $J$ of particles. This is compared for several $(\beta,\kappa)$ combinations. In this subsection, we find the empirical distribution for any $(\beta,\kappa,J)$ by simulating localized CBS $480$ times and, for each simulation, using the positions of $50$ randomly chosen particles at the final time step as samples.
Figure 6: Wasserstein-22 error between the localized CBS samples and the target distribution Eq. 71
@@
As $J$ keeps increasing, the errors in Figure 6 become slightly larger again. This may be caused by a combination of two factors: (i) the Moreau envelope in section 3 approximates $\pi$ by a distribution with longer tails and (ii) when $J$ is small, localized CBS tends to undersample the tails. While the first factor is a source of error present for all $J$, the second factor may partially offset it for small $J$, resulting in an error smaller than that for $J\rightarrow\infty$}.
@@ 7.4 A Darcy flow inverse problem
Consider the spatial domain $\Omega\coloneqq[0,1]^{2}$}. The Darcy flow inverse problem is to recover parameters $u$ of the permeability field $a(x,u)$ from noisy measurements of the pressure field $p(x)$ satisfying
@@
The permeability $a(x,u)$ is modeled as a log-normal random field with mean zero and covariance operator $(-\Delta+\tau^{2})^{-s}$ of $\log a$, where $\Delta$ represents the Laplacian with homogeneous Neumann boundary conditions. This corresponds to a standard normal prior on the parameters $u_{k}$ in the Karhunen–Loève (KL) expansion
$\log a(x,u)=\sum\nolimits_{k\in\mathbb{N}^{2}\setminus\{(0,0)\}}u_{k}\sqrt{\lambda_{k}}\,\phi_{k}(x),$ (73)
with eigenpairs $\lambda_{k}=(\pi^{2}\lVert k\rVert^{2}+\tau^{2})^{-s}$ and $\phi_{k}=c_{k}\cos(\pi k_{1}x_{1})\cos(\pi k_{2}x_{2})$, where $c_{k}=\sqrt{2}$} if $k_{1}k_{2}=0$ and $c_{k}=2$ otherwise (see, e.g., [28]). We choose parameters $(\tau,s)=(3,2)$ and truncate the expansion Eq. 73 after the $d$ terms with largest eigenvalues $\lambda_{k}$. The coefficients $u_{k}$ of these terms are the unknown parameters $u$ with prior $\pi_{\mathrm{prior}}=\mathcal{N}(0,\Gamma_{0})$, with $\Gamma_{0}=I$, in the Bayesian inverse problem Eq. 2. The data $y$ consists of measurements of $p$, taken at $49$ equispaced points and subject to noise known to be distributed according to $\pi_{\mathrm{noise}}=\mathcal{N}(0,\Gamma)$ with $\Gamma=10^{-4}I$. The model $G$ solves Eq. 72 with central finite differences on a grid with stepsize $h=2^{-5}$. An observation is generated by sampling $\{u_{k}\}_{k}$ from the prior, solving Eq. 72 with stepsize $h=2^{-9}$, and perturbing that solution with a random sample from $\pi_{\mathrm{noise}}$. As described in subsection 1.1, this Bayesian inverse problem corresponds to Eq. 1 with
@@
Figure 7: MCMC and localized CBS samples of the first 88 KL terms in the $d$-dimensional Darcy flow problem, with $d\in\{8,16,32\}$. The $d=8$ plot labels the modes in order of decreasing eigenvalue $\lambda_{k}$.
We solve this inverse problem with random-walk MCMC (as a reference solution) and localized CBS. We use $N=400$ timesteps of length ${\Delta\kern-0.8ptt}=0.03$J=1000{J=1000} particles, and a large initial covariance $\Sigma_{0}=9I$ (similarly to [6]). As $\Gamma$ is small, we expect the posterior to be close to Gaussian, which is confirmed by the MCMC reference solution. Hence, we choose moderate parameters $(\beta,\kappa)=(3,0.2)$. Figure 7 compares the MCMC samples to the localized CBS samples. It shows the marginalized distributions for the first 8 modes, for the number of dimensions $d\in\{8,16,32\}$. The modes are successfully located for each $d$. When $d\in\{8,16\}$, the uncertainty around the mode is captured well; when $d=32$, it is slightly overestimated. This matches the argument in Remark 6.1 that scaling localized CBS to high dimensions may prove difficult, although further research is needed.
@@ 8 Conclusions
We have derived localized consensus-based sampling (localized CBS) as an alternative formulation of consensus-based sampling (CBS) obtained from ensemble-preconditioned Langevin dynamics through a sequence of approximations involving the Moreau envelope and a proximal operator. By retaining only the steps that remain valid beyond the Gaussian setting, we obtained a variant closely related to polarized CBS but endowed with full affine invariance. In the mean-field limit, localized CBS is exact for Gaussian targets.
Numerical experiments demonstrate that localized CBS achieves comparable or improved performance relative to polarized CBS and localized ALDI, particularly in non-Gaussian and moderately multimodal problems. The method is fully gradient-free and easily parallelizable.
Several open research directions remain. First, the theoretical understanding of localized CBS—including its non-Gaussian convergence, well-posedness, and mean-field limit—is still incomplete. Second, as discussed in Remark 6.1, the efficiency of localized CBS in high-dimensional settings has not been analyzed and likely depends on problem structure and localization scale. Third, localized CBS may be suitable for sampling from non-differentiable potentials, similarly to other prox-based samplers [34, 45, 58]. Finally, adapting localized CBS to exploit additional problem information—for example, hybrid strategies that use gradient evaluations only for selected particles—may help extend its applicability to more complex sampling tasks.
@@ Appendix A Various technical proofs
Proof of Eqs. 25, 32 and 33.
Consider means $(m,m_{t})$ and covariances $(\Sigma,\Sigma_{t})\hskip-3.0pt$ such that $\pi\sim\mathcal{N}(m,\Sigma)$} and $\rho_{t}\sim\mathcal{N}(m_{t},\Sigma_{t})$. We prove Eq. 25 by computing
@@
which is the mean of the product of two Gaussians, $\mathcal{N}(m,\Sigma/\alpha)$ and $\mathcal{N}(m_{t},\Sigma_{t})$. Similarly to [4], we use the formula found in, e.g., [46, §8.1.8] to compute $\mu_{\alpha}(\rho_{t})$ and, similarly, $\mathcal{C}_{\alpha}(\rho_{t})$:
$\mu_{\alpha}(\rho_{t})=(\alpha\Sigma^{-1}+\Sigma_{t}^{-1})^{-1}(\alpha\Sigma^{-1}m+\Sigma_{t}^{-1}m_{t})\qquad\text{and}\qquad\mathcal{C}_{\alpha}(\rho_{t})=(\alpha\Sigma^{-1}+\Sigma_{t}^{-1})^{-1}.$ (76)
@@
Equations 32a and 32b follow directly from the definitions. Equations 32c, 32d and 33 are derived similarly to $\mu_{\beta,\kappa\mathcal{C}_{\alpha}(\rho_{t})/\beta}(\rho_{t};U_{t})$ above. ∎
@@
In [48, Lemma 3.11], it is shown that with general weights,
$\nabla_{u^{i}}\cdot\widehat{\mathbf{C}}=\omega^{ii}\,(d+1)(U^{i}-\bar{U}^{(i)})+\sum\nolimits_{j=1}^{J}\left(U_{c}^{(i),j}(U_{c}^{(i),j})^{T}-\bar{U}^{(i)}(\bar{U}^{(i)})^{T}\right)\nabla_{u^{i}}\omega^{ij}.\hskip-2.84544pt$ (77)
@@
$\quad=\frac{\omega^{ij}}{\lambda}\Bigl(\bigl[\widehat{\mathcal{C}}^{-1}(U^{j}-U^{i})+\frac{X^{ij}}{2}\bigr]+\bigl[\widehat{\mathcal{C}}^{-1}U_{c}^{(i),i}-\sum\nolimits_{k=1}^{J}\omega^{ik}\frac{X^{ik}}{2}\bigr]\Bigr)+(\delta^{ij}-\omega^{ij})\alpha\omega^{ii}\frac{\nabla_{u^{i}}\widehat{\pi}(U^{i})}{\widehat{\pi}(U^{i})}$
@@
with $(X^{ik})_{\ell}=(U^{i}-U^{k})^{T}\widehat{\mathcal{C}}^{-1}\,(\partial_{u^{i}_{\ell}}\widehat{\mathcal{C}})\,\widehat{\mathcal{C}}^{-1}(U^{i}-U^{k})$. We have $\partial_{u^{i}_{\ell}}\widehat{\mathcal{C}}=J^{-1}\bigl(U_{c}^{i}e_{\ell}^{T}+e_{\ell}(U_{c}^{i})^{T}\bigr)$ and
@@
$B_{1}$ $\coloneqq\sum\nolimits_{j=1}^{J}U_{c}^{(i),j}(U_{c}^{(i),j})^{T}\omega^{ij}\bigl[\widehat{\mathcal{C}}^{-1}U_{c}^{(i),j}\bigr]=\bm{U}_{c}^{(i)}\bigl(\omega^{i}\odot\mathrm{diag}((\bm{U}_{c}^{(i)})^{T}\widehat{\mathcal{C}}^{-1}\bm{U}_{c}^{(i)})\bigr),$